Consider reaction (1) shown below with it's associated equilibrium constant. What is the equilibrium constant for reaction (2)?

Report your answer to THREE significant figures.

   (1)    3 A + 3 B ⇌ 3 C , K = 0.464

   (2)    3 C ⇌ 3 A + 3 B , K = ??
Answer #1

(1). 3A + 3B 3C for this reaction K = 0.464

expression for K = [C]3/[A]3 [B]3 = 0.464. ...... (1)

(2) 3C 3A + 3B.

expression of K for this reaction is

K = [A]3 [B]3/ [C]3 which is reverse of (1)

K = 1/0.464 = 2.16

thus we calculated K of second reaction which is just reverse of (1).